OpenStudy (anonymous):

which of the following is the solution of logx-3^0.001=-3

OpenStudy (anonymous):

a. x=7 b. x=5 c. x=13 d. x=15

OpenStudy (anonymous):

which problem is it? \[\large log(x-3)^{0.001}=-3 \] \[\large logx-3^{0.001}=-3 \]

OpenStudy (anonymous):

thats not even one of the options. i listed them above.

OpenStudy (anonymous):

i'm getting a clarification from you of what the problem is...

OpenStudy (anonymous):

so you were basically just re stating the question?

OpenStudy (anonymous):

yes.. because many times the question isn't really what the asker posted

OpenStudy (anonymous):

oh okay i understand what you did there, well can you help me solve now? please. :)

OpenStudy (anonymous):

i'm guessing the problem is the first one then...

OpenStudy (anonymous):

okay, but i have a test on this tomorrow and i wanna be able to understand everything will you explain to me why you think its the first one?

OpenStudy (anonymous):

no... i'm not talking about the answer yet.. i'm talking about the problem.

OpenStudy (anonymous):

and btw... i will not give answers...

OpenStudy (anonymous):

this isnt even answers, im doing my study guide for tomorrows test this and like 2 other problems are the only ones i dont understand, i put them up here for someone to work out for me so i know how to work similar problems tomorrow, no need to be a jerk.

OpenStudy (anonymous):

he's not being a jerk. he's asking whether your question is the first one or the second out of the two options he gave.

OpenStudy (anonymous):

sorry, i think we're having some kind of communication breakdown here... i was trying to get a clarification of the problem. i was not being a jerk as you stated.. but i'll step aside on this one if that will make you feel better... good luck to you...:)

OpenStudy (anonymous):

thanks nuk64...:)

OpenStudy (anonymous):

which of the following is the solution of logx-3^0.001=-3 <--- the question.

OpenStudy (anonymous):

that's ambiguous! use the equation tool.

OpenStudy (anonymous):

or draw it out.

OpenStudy (lgbasallote):

@sourpatchwatermelon is it \[\log_x -3^{0.001} = -3??\] the problem i mean